### v1

# Detours for Navigating Instructional Videos

Abstract
We introduce the video detours problem for navigating instructional videos. Given a source video and a natural language query asking to alter the how-to video’s current path of execution in a certain way, the goal is to find a related “detour video” that satisfies the requested alteration. To address this challenge, we propose VidDetours, a novel video-language approach that learns to retrieve the targeted temporal segments from a large repository of how-to’s using video-and-text conditioned queries. Furthermore, we devise a language-based pipeline that exploits how-to video narration text to create weakly supervised training data. We demonstrate our idea applied to the domain of how-to cooking videos, where a user can detour from their current recipe to find steps with alternate ingredients, tools, and techniques. Validating on a ground truth annotated dataset of 16K samples, we show our model’s significant improvements over best available methods for video retrieval and question answering, with recall rates exceeding the state of the art by 35%.
1 Introduction
Instructional or “how-to” videos are a compelling medium for people to share and learn new skills. From everyday home fix-it projects, cooking, sports, to aspirational goals like playing piano beautifully, there are so many things that people of all ages and backgrounds want to learn or do a bit better. Indeed online how-to’s are among the top few dominating categories of all content on YouTube, alongside entertainment and music. Advances in computer vision for keystep recognition [44, 88, 21, 6, 87, 54, 55], procedural task understanding [12, 11, 86], and video summarization [58, 5] have the potential to make such content more searchable and accessible.
Figure 1: An example video detour. In the Chicken Quesadillas recipe, the source video $V_{s}$ (top) shows the use of an electric grill at time instant $t_{s}$. A user watching this video does not have a grill and asks a query $\mathcal{Q}$ “how to do this without an electric grill?”. In response, the system identifies a detour video $V_{d}$ and timepoint $T_{d}$ showing a similar recipe but using a heating pan instead of a grill.
However, while today’s how-to content is a vast resource, it is nonetheless disconnected. Human learners access how-to’s in doses of one video at a time, studying the advice and visual demonstrations of one expert at a time. While there are thousands and thousands of videos addressing, for example, “how to repair a bike” or “how to make a samosa”, any given video offers only a single execution of a task using a fixed set of ingredients, tools, approach, and assuming a certain skill level. When those criteria do not align, users face a dilemma whether to improvise, risking “breaking” the final output, or to find and watch another video hoping it better matches their constraints. Manually synthesizing the information across videos is time consuming if not prohibitively taxing.
What if the wealth of knowledge in online instructional videos was not an array of isolated lessons, but instead an interconnected network of information? What would it take to transform a pile of videos into a how-to knowledge base?
Towards this vision, we explore how to intelligently navigate between related how-to videos, conditioned on a natural language query. Suppose a user watching a given video discovers they do not have the desired ingredients, tools, or skill-level. They may ask, “can I do this step without a wrench?” or “I am on a diet, can I skip adding cheese here?” or “how could I prepare the mix from scratch instead of using a pre-made one?” or “is there a simpler way to do the corners?” and so on. Conditioned on the content watched so far in the source video, the goal is to identify a detour video—and a temporal segment within it—that would allow the user to continue their task with the adjustment specified by their language query, then return to the original source video and complete execution. See Figure 1.
At the core this requires new technical advances in multimodal video understanding. Standard text-to-video retrieval models [54, 73, 55, 17, 85, 22, 49] are insufficient because the query text alone may not reveal enough details about the task (e.g., “can I do this step without an electric grill?”) Similarly, existing video localization [83, 81, 78, 46] and question answering [62, 77, 40, 36, 76, 37] methods do not consider the viewing history of the source video, which is essential to properly identify a detour. For example, answering “how to do this step without a wrench?” requires the model to understand which steps are already done and which part of the target detour video shows the same effect without using a wrench.
We introduce VidDetours: a video-language model that benchmarks this new problem. Our approach formulates the video navigation task in two parts: (a) retrieval of the detour video (b) temporal localization of the relevant portion of the detour video, both conditioned on the source video and text query. Building on ideas from the video retrieval [55, 73, 49, 89] and localization [81, 83, 57] literature, we develop an architecture and training objective that accounts for all the essential components of this task: gauging the relatedness of any two instructional videos; capturing the interchangeability of their component steps; and interactively indexing into the alternatives with language.
Since there are no existing datasets labeled for video detours, we devise a framework leveraging large language models (LLMs) to generate weakly-supervised training data. Using HowTo100M [54], a large-scale instructional video dataset of in-the-wild how-to’s accompanied by the transcribed speech of the narrator, we automatically generate plausible user queries at targeted timepoints in training source videos together with their detour counterparts in closely related videos. This procedure makes it possible to obtain ample effective training data without manual annotations. To rigorously evaluate the video detours, we introduce a gold-standard test set comprised of manually labeled data from 4K full-length videos and 16K human-generated questions.
In extensive experiments, we validate our model and illustrate the promise of the novel task. VidDetours strongly outperforms state-of-the-art video-language and video retrieval methods. We will release our training and test annotations to establish a formal benchmark for navigating instructional videos.
In short, ours is the first work to investigate personalized query-based navigation of instructional videos. Our main contributions are the innovative task definition, our video-language model to address it, and the high quality eval set and benchmark. These results help pave the way towards an interconnected how-to video knowledge base that would transcend the expertise of any one teacher, weaving together the myriad of steps, tips, and strategies available in existing large-scale video content.
2 Related Work
Figure 2: Overview of the detours dataset ($\mathcal{D}_{D}^{tr}$) curation. Given unlabeled instructional videos for training (we use HowTo100M [54]), we first input their narrations with timestamps to a language model (LLAMA2 [66]) to obtain summaries of their steps. Next, we automatically select pairs of similar summaries along with their timestamps and use a language model to generate weakly-supervised detours annotation tuples $(V_{s},t_{s},\mathcal{Q},V_{d},T_{d})$. As an example, the source video here uses smooth peanut butter. A possible detour question is “can I use chunky peanut butter here?” and the window at $T_{d}$ in the detour video (top right, orange) shows the use of crunchy peanut butter.
Learning from instructional videos. Several recent video datasets like HowTo100M [54], COIN [65], CrossTask [90] are based on instructional videos and have enabled research in procedure planning [12, 11, 86], task graph learning [6, 87, 21] and alignment detection [5, 31]. Availability of large-scale instructional videos on the internet also facilitates video representation learning that is used for action recognition [29, 42, 39, 71, 23], action anticipation [28, 24, 51, 1, 26], and object detection [3, 10]. All the prior works either focus on short-term representations [55, 73, 49, 15, 14] or video-level understanding [6, 21]. To our knowledge, ours is the first work to establish a means to navigate across instructional videos, which is essential for holistic task understanding and optimal skill learning.
Vision and language learning. Videos often also contain text—whether converted from narrations through automatic speech recognition (ASR) [54, 65, 90] or manually annotated [30]. Using both text and video for representation learning [55, 73, 43, 4, 15, 14] helps in multi-modal tasks like retrieval [17, 85, 22, 73, 49], localization [62, 73, 78, 90, 80, 81, 46, 74], captioning [33, 59, 41, 79, 69], question answering [77, 40, 36, 76, 37], and episodic memory queries [30, 43, 62]. Most of these tasks focus on images or clip-level understanding, typically a few seconds long. Recent work [4, 7, 25, 53, 60, 30, 62] extends this further for video-level understanding spanning minutes. None of the existing methods and benchmarks answer text queries by navigating between long videos, as we propose.
Interactive retrieval. Dialog-based retrieval has been previously studied in the context of fashion image retrieval [16, 72, 32] and conversation based e-commerce shopping [64, 84], where a user wants a specific product and gives feedback on successive retrievals. In Visual Dialog [18, 19, 13, 34, 56], an agent is given an image and its caption and has to answer questions about the image e.g. “what color is the mug?”. Recent work in composed image/video retrieval [67, 68, 75, 9, 8, 47, 35] use an image/clip with a modification text to retrieve an improved version, e.g. a fountain image with text “at night” retrieves a clip of the fountain at night.
All the previous work focuses on improving video or image retrieval through dialog and the inputs are image or short-duration video. In contrast, we focus on action demonstrations where the prompt can be about ingredients, tools, or even step executions, e.g., “how to prepare the mixture instead of using a pre-made mix?” which is crucial for a holistic task-level understanding of many actions and dependencies amongst them. Furthermore, our setup considers full instructional videos that typically span several minutes, as opposed to static images and short clips.
3 Approach
In this section, we first define our detour task formulation (Sec. 3.1). Next, we detail the dataset collection process (Sec. 3.2) and our model architecture (Sec. 3.3). Finally, we discuss implementation and training details (Sec. 3.4).
3.1 Video detour task formulation
We define a video detour as a mapping from a source video $V_{s}$ at timestamp $t_{s}$ to a response segment $T_{d}=(t_{d}^{s},t_{d}^{e})$ in a detour video $V_{d}$, based on a query text $\mathcal{Q}$. This is illustrated in Fig. 1, where after watching the source video for some time (purple bar, top panel), a user issues a query “how to do this without an electric grill?”, for which the response is a segment in a different video showing the step in a pan instead of a grill (green bar, bottom panel). By construction, $V_{s}$ and $V_{d}$ are related demonstrations from the same high-level task, that differ slightly in their demonstration (e.g., two videos demonstrating how to make chicken quesadillas).
Formally, we cast this as a video segment retrieval task conditioned on both a source video and a query text. The goal is to find functions $\mathcal{F}_{R}$ and $\mathcal{F}_{L}$ such that
$V_{d}$ $=\operatorname*{\arg\!\max}_{V_{i}\in\mathcal{D}}\mathcal{F}_{R}(V_{i}|V_{s}[1:t_{s}],\mathcal{Q})$ (1)
$T_{d}$ $=\operatorname*{\arg\!\max}_{T_{i}}\mathcal{F}_{L}(T_{i}|V_{s}[1:t_{s}],\mathcal{Q},V_{d})$ (2)
where $V_{s}[1:t_{s}]$ refers to a video watched from the beginning, until time $t_{s}$, and $T_{i}$ refers to a temporal window (start and end time) in video $V_{d}$.
Here, $\mathcal{F}_{R}$ is a retrieval mapping that finds the correct full instructional video, typically minutes long, given the source video segment and the text query, while $\mathcal{F}_{L}$ is the localization function that finds the start and end time in the detour video.
3.2 Detour dataset generation
Our goal is to learn retrieval and localization functions to find the correct detour segment given a source video and a query, for which we require a training dataset with tuples of the form $(V_{s},t_{s},\mathcal{Q},V_{d},T_{d})$. The detour queries can focus on any aspect of the demonstration of the recipe: ingredients (e.g., “can I add eggs here?”), tools, (e.g., “how to do this without an blender”) or steps (e.g., “how do I serve this on a plate?”). Existing procedural video datasets only offer a subset of the required information [54, 90, 65, 2]: they provide narrations or keystep labels, but are missing interrelations between different procedural demonstrations and thus cannot be used for detours training directly. Moreover, collecting detour annotations at a large-scale may be impractical due to the amount of time required for annotators to watch and parse long detour videos. To address this, we propose an approach to automatically create a training dataset $\mathcal{D}_{D}^{tr}$ for our task using how-to video narrations and language models. Subsequently, for rigorous testing, we also manually collect ground truth test data $\mathcal{D}_{D}^{te}$ from human annotators. Each of the dataset generation processes is detailed below.
Weakly-supervised training set $\mathcal{D}_{D}^{tr}$. We start with a dataset of unlabeled instructional videos: $\mathcal{D}=\{(V_{i},N_{i})\}_{i=1}^{|\mathcal{D}|}$ containing narrations $N_{i}$ in addition to the videos $V_{i}$. A narration is the spoken component of the how-to video, where the expert describes their actions (“now we mix for 3 minutes”) and gives other color commentary (“oh it looks great!”). We concentrate on the broad domain of cooking due to its prominence in current instructional video datasets ($\sim$370K videos in HowTo100M [54]), well-structured recipes, and strong interconnection between different instances.
We use the narrations to generate labels for our training dataset $\mathcal{D}^{tr}_{D}$. Despite being noisy, narrations have been used successively for weakly-supervised training labels for video-language pretraining [54, 55, 73, 49] and keystep recognition [6, 44, 87]. Here we explore their utility for mining candidate detour pairs. We do this in two stages. See Figure 2. First, we generate text summaries for the key steps in each video. Specifically, we prompt LLAMA 2 [66], a recent open source large language model, to summarize the instructional videos using the timestamped narrations $N_{i}$. The prompt is of the form “Given the following narrations from a YouTube video, what recipe is this, and summarize each step along with its timestamps…” The exact text prompt is given in the Supp. along with example outputs. We obtain the summary of video $V_{i}$ as a tuple of the step start time, end time, and text description. This process yields an intermediate large-scale summary dataset, see Figure 2 (second panel).
Next, we generate detour queries and time windows given a pair of summaries. For this, we identify video pairs that share an activity (and therefore have similar summaries), as unrelated pairs are unlikely to yield a meaningful detour query $\mathcal{Q}$. Specifically, we sort summary pairs by cosine similarity of their MPNet [63] sentence embeddings, discarding dissimilar pairs (score $<0.75$). With a video pair in hand, we design a prompt for the LLM to generate detour queries and time windows $(t_{s},\mathcal{Q},T_{d})$ of roughly the form “Given video summaries with timestamps, suppose a person is watching video A, identify a text prompt that a user might issue to take a detour and watch video B, along with the detour timestamps? Some examples of detours …”. See Supp. for full prompt details. Note that a source video can have multiple valid queries and matching detour videos, which our generation strategy allows.
Figure 3: Visualization of most frequent bigrams of the queries in the manually annotated test set. We see that most of the queries have little or no context about the current recipe and the step being executed, e.g. “how do I plate this differently?” or “can I do this step using a spatula?”—emphasizing the need for source video context.
The entire process yields a pair of videos and the detour annotations $(V_{s},t_{s},\mathcal{Q},V_{d},T_{d})$ (Figure 2, right panel). More examples are in Supp. While $\mathcal{D}_{D}^{tr}$ will naturally have some noise due to language model errors and misaligned or non-visual narrations [31, 5], we find them quite reasonable from manual inspection of a subset of the data. More importantly, they are effective for training a detour model, as our experiments with manually labeled data will show.
Manually collected testing set $\mathcal{D}_{D}^{te}$. While the weakly-supervised data is sufficient for training, for reliable evaluation of our trained models and baselines, we manually collect ground truth test data. Similar to $\mathcal{D}_{D}^{tr}$, we identify a pair of similar videos, and ask the professional annotators to watch the videos completely, and then annotate $(t_{s},\mathcal{Q},T_{d})$. Annotators are allowed to reject pairs for which detours cannot be constructed (e.g., if they are too dissimilar). Since there can be multiple detours possible for a given pair of videos, we ask the annotators to identify at least three detours. We ensure that the videos in the train and test set are disjoint. The annotation process results in a high quality, benchmark test set for our new task.
Figure 3 shows the diversity in queries that annotators provide. We see a variety of questions arise about substitutions of ingredients, tools, and steps. Further, we see contextual queries, e.g. “can I skip adding butter?” that do not reveal much information about the recipe and the current step, underscoring the need to reference the source video when localizing a detour. The dataset statistics and splits are given in Sec. 3.4.
Figure 4: Overview of the proposed approach. The source video and the candidate detour video are converted to visual tokens by passing through a video encoder $f_{v}$, followed by a visual mapper $f_{vm}$. We obtain a similar text token using standard tokenizer $f_{n}$. The processed tokens are then passed through a multimodal sequence encoder $\mathcal{L}_{se}$ to obtain output features. Finally, we have specific task heads for detour video retrieval and detour window localization.
3.3 Detour retrieval and localization modules
Next, we describe our training framework to learn the mapping functions $\mathcal{F}_{R}$ and $\mathcal{F}_{L}$ using our generated detour dataset. Our objective is to design a multimodal (video and text) architecture that fuses the source video context with the language query enabling $\mathcal{F}_{R}$ and $\mathcal{F}_{L}$ to utilize both the viewing history and the query. As we will see in the experiments, this idea is more effective than late fusion of video and query features [70, 61, 7]. For this, we leverage the reasoning capabilities of large language models (LLM). In short, we encode both videos and the detour query as a sequence of tokens to pass to the LLM, which aggregates multimodal information from the inputs in its output encoding. These encodings are finally used to retrieve detour videos and detour segments using task specific heads. Our overall framework is shown in Fig. 4 and each component is described in detail below.
Encoding videos and detour queries as sequences of tokens.
We begin by encoding the videos as a sequence of tokens compatible with the LLM. We encode each video as spatio-temporal features using a video encoder $f_{v}$ (e.g., InternVideo [70]), extracted at one feature per second [73, 6] Next, we use a low-parameter visual mapper $f_{vm}$ [45, 38, 27, 50, 82] to convert the video features into visual tokens v $=~{}f_{vm}(f_{v}(V))$, that are in the same embedding space as the text tokens, making them compatible with language encoders.
[…]
LLMs as multimodal sequence encoders.
Next, we use a LLAMA2 model as our multimodal sequence encoder ($\mathcal{L}_{se})$. LLMs are an ideal choice here, given our goal of encoding dialogue-driven detours — the idea that a user is watching the source video and pauses it to ask a query, followed by a response from the LLM. Specifically, we append source video tokens (until time $t_{s}$) with the query tokens $\text{{v}}_{s}[1:t_{s}]~{}|~{}\text{{n}}_{\mathcal{Q}}$, followed by candidate response video tokens. The multimodal encoder captures the cross-modal interactions between the source video, query context and the candidate video, resulting in an encoded output for the candidate video $O_{i}=\mathcal{L}_{se}\left(\text{{v}}_{s}[1:t_{s}]~{}|~{}\text{{n}}_{\mathcal{Q}}~{}|~{}\text{{v}}_{i}\right)$.
Detour retrieval and localization heads
Finally, we use the updated multimodal encodings $O_{i}$ to score detour video candidates and segments using two task heads $f_{R}$ and $f_{L}$. $f_{R}$ is a classifier that scores the relevancy of candidate video $V_{i}$ given the viewing context and the user query, while $f_{L}$ is a classifier that identifies the highest score time segment $T_{i}$ inside the correct detour video $V_{d}$, given $O_{d}$. These two heads are trained separately for each task.
For our video retrieval network $\mathcal{F}_{R}$, we minimize the binary cross entropy loss $f_{BCE}$ between the prediction and the ground truth label, i.e.
$\operatorname*{\min}_{V_{i}\in\mathcal{D}}f_{BCE}(\mathcal{F}_{R}(V_{i}|V_{s}[1:t_{s}],\mathcal{Q}),\mathbbm{1}(V_{i},V_{d})).$
We assign a positive label to the correct detour video $V_{d}$ and negative to other videos sampled from the dataset, i.e. $V_{i}\neq V_{d}$. In particular, for every correct training instance, we randomly sample an incorrect video from which to curate a negative sample—either from the same task (hard negatives) or other tasks.
For the localization network $\mathcal{F}_{L}$, the training objective is:
$\operatorname*{\min}_{T_{i}=[t_{i}^{s},t_{i}^{e}]}\frac{1}{2}\left[f_{CE}(t_{i}^{s},t_{d}^{s})+f_{CE}(t_{i}^{e},t_{d}^{e})\right]$
where $f_{CE}$ is the cross-entropy loss. Similar to the video-language grounding model VSLNet [81], this objective minimizes the error in the distribution of the start and end times across the video. At inference, we find the candidate video $V_{d}$ and time duration $T_{d}$ that maximizes the scores from $\mathcal{F}_{R}$ and $\mathcal{F}_{L}$, respectively (see Sec. 3.1). Note that there may be multiple plausible detour videos for a given source and query (e.g., multiple videos can use a heating pan instead of an electric grill), for which our scoring-based approach ensures other related (and valid) pairings can also score highly.
Method
R@5
R@10
R@50
MedR $\downarrow$
Text-only 3.9 8.7 14.0 512
CLIP [61]
7.9
11.8
25.2
342
CLIP-Hitchhiker [7]
8.4
12.3
25.6
336
InternVideo [70]
9.7
13.2
27.2
313
CoVR [67]
4.3
9.2
15.3
473
Ours
17.6 27.8 62.4 30
Ours w/o hard-negatives
16.5
24.9
56.3
55
Ours w/ parser
13.9
21.6
50.0
81
Method
R@1, IoU=0.3
R@1, IoU=0.5
R@1, IoU=0.7
Mean R@1
Text-only 5.2 2.7 0.6 4.2
2D-TAN [83]
10.3
4.2
1.5
8.6
VSLNet [81]
11.8
5.8
1.7
9.4
UMT [46]
12.0
6.1
1.6
9.4
STALE [57]
12.1
6.1
1.7
9.6
Ours
16.7 7.7 2.8 12.8
Ours w/ parser
13.4
7.0
2.5
11.6
Table 1: Results for detour video retrieval (left) and detour window localization (right) tasks. Our method outperforms all prior methods and baselines by a significant margin.
3.4 Implementation details
Dataset and statistics. Our training and test sets are both derived from HowTo100M [54] (average length 6.5 mins). We consider cooking tasks (i.e., recipes) containing $370K$ videos. We use the weakly-supervised automatically curated dataset $\mathcal{D}_{D}^{tr}$ for training and validation. Following the steps described in Sec. 3.2, we obtain $586,603$ training and $18,308$ validation detour annotation tuples $(V_{s},t_{s},\mathcal{Q},V_{d},T_{d})$.
The manual annotation results in a large-scale test dataset $\mathcal{D}_{D}^{te}$ containing $16,207$ detour instances $(V_{s},t_{s},\mathcal{Q},V_{d},T_{d})$. The test set is based on $3,873$ unique videos across $1,080$ recipes/tasks, e.g. “how to make chicken quesadillas” is one task and has multiple video instances. We curate the test set so that there are $834$ common tasks ($14,450/16,207$ annotations) with the training data $\mathcal{D}_{D}^{tr}$ having videos from those tasks. There are $246$ additional novel recipes ($1,757/16,307$ annotations) that do not contribute to the training set. For retrieval evaluation, the detour candidates are all videos in the dataset, i.e. $3,873$ candidates per detour annotation. No video exists in both the training and testing split.
Network architecture. We use InternVideo [70] as the video feature extractor $f_{v}$. The features are extracted at one feature per second, following [73, 31]. $f_{v}$ is frozen and hence, the features are pre-extracted and saved offline for faster I/O. Next, $f_{vm}$ is a trainable linear layer, inspired by [45]. $\mathcal{L}_{se}$ is a LLAMA2-13B-chat [66] language model. We try both variants of keeping $\mathcal{L}_{se}$ frozen and trainable and performance is better if $\mathcal{L}_{se}$ is trainable. Lastly, $f_{n}$ is the LLAMA-2 tokenizer. The retrieval head $f_{R}$ is a transformer classifier and we take the CLS token of the output followed by a linear layer to output the score. Finally, the localization head $f_{L}$ is the VSLNet [81] architecture. We remove the tokenizer from VSLNet since the text input is already processed.
Training Parameters. We train both networks on 8 nodes with 8 NVIDIA A100 GPUs for 5 epochs. The training time is 8 hours. We use AdamW [48] optimizer with learning rate $3\times 10^{-5}$ and batch size of 16 per device. The transformer classifier $f_{R}$ uses an input dimension of 4096 (consistent with the LLAMA2 output dimension), 4 heads, 4 layers and 1024 dimensional feed-forward network. All other parameters are defaults of the respective models.
4 Experiments
Figure 5: Visualization of our model’s predictions for the detour video retrieval and window localization tasks. For all the examples, our method ranks the detour video at top-1. Furthermore, in the detour window localization, our method is able to predict the detour window with a high overlap ($>0.7$ IoU). Our model is able to correctly use source video context and nuanced queries like “how to cut it in star shape?” (Example 6) and “can I skip the onions?” (Example 3).
We show the results for the detour video retrieval (Sec. 4.1 and detour localization (Sec. 4.2) subtasks.
4.1 Detour video retrieval
We benchmark models on finding the correct detour video given the source video and the query, from amongst all videos in the test set (3,873 videos).
[…]
  • with $V_{s},\mathcal{Q}$: where $\psi(V_{s},\mathcal{Q})=1/2[\phi(V_{s})+\phi(\mathcal{Q})]$.
  • with $V_{s}$: where $\psi(V_{s},\mathcal{Q})=\phi(V_{s})$.
  • with $\mathcal{Q}$: where $\psi(V_{s},\mathcal{Q})=\phi(\mathcal{Q})$.
These three variants capture whether individual embeddings (or simple combinations of them) are sufficient for detour retrieval.
  • Text-only computes the similarity between summary and query embeddings, ignoring visual cues. This baseline evaluates the impact of text-bias in the automatically curated dataset.
  • CLIP [61], InternVideo [70] are state-of-the-art vision-language models used extensively for multi-modal tasks. Since they take short video clips as input, the video representation is the average of all short-term features.
  • CLIP-Hitchhiker [7] is similar to CLIP [61] but uses a weighted average of frame features, instead of uniformly averaging for $\phi(V_{i})$. The weights are the similarity score between the frame and query features (eq. 1 in [7]). Note that the source-only variant is not evaluated as query features are needed to compute the weights.
  • CoVR [67] is a state-of-the-art method for composed video retrieval. Following CoVR, we sample a frame from the source video and use the user query as the modification text to obtain detour retrieval.
Ablations. Similar to how we evaluate the baselines with different inputs, we compare the performance of our method (i.e. “with $V_{s},\mathcal{Q}$”) with ablations — “with $V_{s}$” and “with $\mathcal{Q}$”. These baselines capture the performance when no query is provided and when source video $V_{s}$ is not given to the model, respectively. Next, recall that we use a language model to generate timestamps for both the summaries and the detour timestamps (see Fig. 2 and Sec. 3.2) that can be noisy. Instead, for this ablation, we parse the summaries and queries and assign timestamps based on the similarity score between the narrations and the summaries. The language model does not handle timestamps for this “Ours w/ parser” baseline. Finally, we evaluate the retrieval performance without sampling negatives from the same task i.e. no hard negatives in “Ours w/o hard negatives” baseline.
Metrics. Following standard retrieval evaluation [54, 55, 73, 52], we report recall@$k$ for $k\in[5,10,50]$ and the median rank. Recall@$k$ ranges in $[0,1]$, higher the better. The median rank ranges between $1$ and the size of the retrieval candidate set (i.e. $3,873$), lower the better.
Results. We report the results with both inputs, i.e. “with $V_{s},\mathcal{Q}$” in Table 1 (left) for all baselines on all metrics. We also report medR for different inputs in Table 2 (left). Our method outperforms all the baselines by a significant margin. Our performance gain is 7.9% w.r.t. InternVideo [70] at R@5 and the gap further increases to 35.2% for R@50. Across all the different input combinations, InternVideo [70] “with $\mathcal{Q}$” is the second best and attains a medR of 138 — much lower than our medR of 30 (Tab. 2).
We attribute our large gain to three crucial factors — (a) thanks to its tokens and long sequence length in the LLM, our model is capable of long video understanding, unlike CoVR [67], which samples a few frames from the video, (b) appropriately using the source video context $V_{d}$, unlike CLIP [61], CLIP-Hitchhiker [7], and InternVideo [70] since our model is more than a standard text-to-video retrieval model and (c) conditioning the retrieval on the query text.
Fig. 5 shows example detours inferred by our model. In all the cases, our retrieval model ranks the correct detour video at top-1. It is evident that detour video retrieval is more challenging than conventional text-to-video-retrieval because the detour query can have missing context, e.g. “how do I add garnish after pouring?” does not reveal the video task; the previous video context is needed for the correct retrieval. See Supp. for more visualizations, comparisons and some failure cases. The performance trend is similar for common and novel tasks of the dataset, reinforcing that training on a sufficiently large dataset with many recipes enables good generalization. We show the performance of each test subset in the Supp.
4.2 Detour window localization
Method
$V_{s}$
$\mathcal{Q}$
MedR $\downarrow$
CLIP [61]
314
191
342
CLIP-Hitch. [7]
186
336
InternVideo [70]
150
138
313
CoVR [67]
388
401
473
Ours
128
116
30
Method
$V_{s}$
$\mathcal{Q}$
R@1
2D-TAN [83]
5.5
8.0
8.6
VSLNet [81]
6.1
8.5
9.4
UMT [46]
6.5
8.7
9.4
STALE [57]
6.9
8.8
9.6
Ours
8.9
11.2
12.8
Table 2: Comparison of our method with prior methods at different input combinations for detour video retrieval (left) and detour window localization (right). Our method outperforms all the prior works for all input combinations. See Supp. for all metrics.
Next, we show results on detours window localization where the task is to determine the correct window given the query, the source video, and the ground-truth detour video.
Baselines. While there are no existing works that use previous minutes-long viewing history along with the query to perform temporal localization, we use state-of-the-art localization works as our baselines and also strengthen them by providing source video context, as described below.
  • Text-only: Similar to the detours video retrieval, we have a text-only baseline to evaluate the text-bias in the automatically curated train set.
  • 2D-TAN [83], VSLNet [81], UMT [46]: All these prior models aim to localize text in videos. Thus, we also propose variants of these methods at different input combinations, same as the detour retrieval task. To provide video context, we use a visual mapper $f_{vm}$ similar to our proposed approach—thus providing a late fusion of video $V_{s}$ context. These visual tokens are prepended to the text token, same as our token sequence (see Fig. 4). This enhancement enables us to evaluate “with $V_{s}$” and “with $V_{s}$, $\mathcal{Q}$” in addition to the standard “with $\mathcal{Q}$
  • STALE [57]: This is zero-shot temporal detection method uses vision-language prompting. Same as above, we evaluate this baseline at three combination on inputs.
Metrics. Following [30, 81, 83], we report recall@1 for IoU thresholds in $[0.3,0.5,0.7]$. We also report the recall@$1$ at the average IoU. All the recall metrics range in $[0,1]$, higher the better.
Results. Tab. 1 shows the results with maximum supervision (“with $V_{s}$, $\mathcal{Q}$”) and Tab. 2 (right) compares against different input combinations. We use the same set of ablations as the detour retrieval task, see Sec. 4.1. We clearly outperform all the baselines and ablations by a clear margin. Our mean recall@1 is 3.2% more than the second best performing method STALE [57]. The same trend is true for all IoU thresholds, with higher threshold having lower recall, as expected. Again, the LLM-based parser is better for obtaining the timestamps.
Our gains can be attributed to our model design involving early fusion of previous viewing context and the query features. The detour queries have less context than in a typical text localization task. For example, “can I skip adding salt here?” requires a model to first understand the step being done in the source video, followed by interpreting the query to localize a similar step without salt. Existing methods are incapable of capturing this multimodal dependency, even if we strengthen them with late fusion. We also see that our mean recall@$1$ performance is better than all other methods at all input combinations, including ablations.
Fig.  5 shows example detour localization (right) for our method. We see that our method is able to use the source video context and the user query to localize the detour window. For example, when a user asks “can I do this step with a steel plank saver”, our model correctly localizes the process of using a steel plank saver to put the salmon of the grill in the target video. See Supp. for more visualizations and comparisons.
5 Conclusion
We propose a novel task of finding detours for navigating instructional videos. Building on video and language modeling, we develop a weakly-supervised training dataset and a novel method to train a detours network. Our results show how existing similarity-based retrieval and localization methods are insufficient to address this problem. The dataset will be released to the community to support research in navigating instructional videos.
Supplementary Material
1 Supplementary video
We attach a supplementary video containing details about high-level idea, overview of the problem, detour dataset visualizations and some result visualizations.
2 Dataset visualization
We attach visualization pages that show the outputs from LLM for (a) summary generation and (b) weakly-supervised detour training data. Both these visualizations contain samples that are automatically generated using narrations with LLAMA 2. Most of the training samples are valid detour prompts with correct time windows; with some noise due to imperfections in narrations as a supervision and due to errors in LLAMA 2 generations. Finally, we also have a visualization that shows the samples from (c) manually collected testing data. These visualizations show the good quality manual annotations that we have collected.
3 Detour dataset generation details
In this section, we present additional details about the detour dataset generation (Sec. 3.2 in the main paper) process. We discuss the input prompt used to generate weakly-supervised detour annotations in $\mathcal{D}_{D}^{tr}$ and resulting sample outputs (including failure cases). Next, we detail the manually annotated dataset $\mathcal{D}_{D}^{te}$ collection details and sample visualizations.
3.1 Weakly-supervised training set.
The first step involves generating summaries from given narrations $N_{i}$. The narrations are obtained using ASR from narrated videos and we use the sentencified version, provided in [31]. We use the narrations along with the timestamp and provide the following prompt to LLAMA 2 [66]:
{myframe}
System: Help summarize the steps of this YouTube recipes whose narrations with timestamps are given. Timestamp is given in HH:MM:ss.
User: Given the narrations of a YouTube video, tell the recipe being made in this YouTube video and list down the steps and start and end timestamps in the video. Answer in this format: ‘Recipe: Name of the recipe and brief detail Step 1: [HH:MM:ss - HH:MM:ss] description of the step Step 2: [HH:MM:ss - HH:MM:ss] description of the step and so on’. Here are narrations with timestamps in HH:MM:ss format: $<$insert$>$
Figure 1: Weakly-supervised summaries generated using narrations with LLAMA 2 [66]. While majority of the outputs contains step details and timestamps in the desired format, a few outputs are incorrect (bottom).
where $<$insert$>$ is replaced by narrations with timestamps. Some sample outputs are shown in Fig. 1 along with a row of failure cases (bottom). We create an automated parser that extracts steps as a tuple of timestamps and text description. Many of the videos do not contain meaningful narrations or have no narrations, and hence the outputs from these prompts do not fit into the desired output format. They are rejected automatically by the parser. There are rare instances where even though the narrations are meaningful, the outputs are incorrect, e.g. garbage output or no output (bottom right, Fig. 1). Finally, there is a small fraction of outputs ($\sim$3%) where the timestamps are incorrect or missing altogether (bottom left, Fig. 1). To mitigate this, we make sure steps’ coverage is at least 80% of the duration of the video. This process results in a high-quality text summary dataset. The overall process results in a summary dataset of 187K samples. Please also see the attached summary visualization page. It contains parsed summaries and we can observe the good quality summary generations using LLAMA 2 with only few failure cases.
Finally, we input two similar summaries into LLAMA 2 [66] and generate detour instances. The process to filter similar summaries is detailed in Sec. 3.2 (main paper). For every pair of similar summaries, we use the following prompt:
{myframe}
System: Help understand why a user would pause watching one video and take a detour to another cooking video.
User: There are two cooking videos A and B. The steps of the recipe along with timestamps in HH:MM:ss format is given. Suppose a person is watching video A, can you tell me what the user would prompt to take a detour and watch video B? The answer can be some extra/missing ingredients, tools or procedural step. Some examples of such queries can be ‘How to do this step without adding yeast?’, ‘Can I add chilli powder here?’, ‘Can I do this step without blender?’, ‘Can you give a video that shows other way to roll a sushi?’ and so on. Also, tell the time when the user would stop watching Video A and the time range in Video B and answers the user query. Answer in this format: ‘Detour time in Video A: HH:MM:ss, Detour time window in Video B: [HH:MM:ss - HH:MM:ss], Detour text prompt: One sentence question a user would prompt to take a detour’. Here are the reciped: Video A: $<$insert$>$ and Video B: $<$insert$>$
Figure 2: Weakly-supervised detour annotation sample for training and validation. It also contains a row of failure cases with reasons. Please also see the attached visualization for more visualizations.
Figure 3: Annotation interface for manual test dataset collection. The interface reiterates important details in addition to a separate document (top). There is a scratch for the annotators to take notes while watching the two videos on the right. Finally, each instance of annotation contains a detour start time from the source video, a detour text prompt and finally a detour time window in the target video. The interface supports up to 15 annotations but only three is required.
where we insert source and detour video narrations in $<$insert$>$, respectively. Same as above, we create an automated parser to convert the text outputs into dataset tuples. A small fraction of outputs cannot be parsed by the automated parser due to incorrect output format by LLAMA 2. We ignore these instances since they are small in number in comparison to the successful parse. Fig. 2 shows some output samples along with failure cases. Please also see the attached visualization page for training data samples that contains automatically annotated valid detour annotations and some failure cases. We manually verify a subset of the generations and observe good quality. Furthermore, we observe a strong correlation coefficient of $>0.85$ between validation set (created automatically using narrations) and the manually collected test set across all training runs for both detour video retrieval and detour window localization task.
This automatically generated data is used for training only—never for ground truth evaluation of any model.
3.2 Manually collected testing set
We hire 24 professional annotators for manually generating video detour instances. All of them are trained for a few days on what constitutes a detour (along with examples), how to mark the time instances $t_{a}$ and detour window $T_{d}$ and what types of samples to reject. The training was followed by a pilot collection to evaluate their understanding. Finally, they annotate using a designed interface shown in Fig. 3. We randomly sample a subset and manually verify the annotations for quality control.
Figure 4: Manually collected detour annotation for testing. Please also see the attached visualization for more samples with videos that showcases the good quality annotations that we collected.
The resulting dataset consists of 3.9K source-detour video pairs resulting in $16,207$ samples. Due to our annotator trainings and quality control, the resulting dataset is of high-quality. Fig. 4 shows representative examples. Please also see the attached visualization of test data samples that shows the high-quality samples. These manually created detours are used for evaluation.
4 Experimental results expanded
We expand on to the results in Sec. 4 (main paper) and show the generalizability of our method (Sec. 4.1) and performance at different input combinations (Sec. 4.2).
4.1 Generalizability to novel tasks
As discussed in Sec. 3.4, we have two splits of the test data — common tasks containing video pairs from most frequent recipes of HowTo100M [54], and novel tasks consisting of video pairs from least frequent recipes of the dataset. We do not include any video pairs from novel tasks in the training set to evaluate the generalizability of our method.
Results. Tab. 1 contains the performance split for each testing subset for both detour video retrieval and detour window localization tasks. In the main paper, we showed only the performance on overall dataset for conciseness. We see that our method achieves significant gains over the baselines for both the tasks. Moreover, the performance drop in novel tasks is minimal compared to the gain. This result shows that the learned model is able to generalize to detour in newer recipes without being explicitly trained on them. We attribute this effect to the strong interconnected nature of demonstrations in instructional videos, cooking in particular. Some recipes like making pancake and making crepe will only differ at some steps and detour learned with making pancake should transfer to making crepe.
Method
Common Task MedR $\downarrow$
Novel Task MedR $\downarrow$
MedR $\downarrow$
Text-only 508 524 512
CLIP [61]
348
310
342
CLIP-Hitchhiker [7]
339
305
336
InternVideo [70]
315
296
313
CoVR [67]
464
485
473
Ours
29 35 30
Ours w/o hard-negatives
49
63
55
Ours w/ parser
76
88
81
Method
Common Task Mean R@1
Novel Task Mean R@1
Mean R@1
Text-only 4.0 4.5 4.2
2D-TAN [83]
8.9
8.2
8.6
VSLNet [81]
9.2
9.8
9.4
UMT [46]
9.6
9.3
9.4
STALE [57]
9.7
9.5
9.6
Ours
13.3 12.3 12.8
Ours w/ parser
12.0
11.3
11.6
Table 1: Results for detour video retrieval (left) and detour window localization (right) tasks on common task and novel task splits. Our method outperforms all prior methods and baselines by a significant margin even on novel tasks.
Method
$V_{s}$
$\mathcal{Q}$
R@5
R@10
R@50
MedR $\downarrow$
CLIP [61]
9.6
13.2
26.9
314
11.2
16.4
32.0
191
7.9
11.8
25.2
342
CLIP-Hitch. [7]
11.3
17.7
33.2
186
8.4
12.3
25.6
336
InternVideo [70]
11.2
17.0
31.8
150
13.1
19.2
37.2
138
9.7
13.2
27.2
313
CoVR [67]
4.9
10.1
15.9
388
4.1
10.0
15.6
401
4.3
9.2
15.3
473
Ours
6.1
11.0
32.3
128
6.1
10.8
32.6
116
17.6 27.8 62.4 30
Method
$V_{s}$
$\mathcal{Q}$
R@1, IoU=0.3
R@1, IoU=0.5
R@1, IoU=0.7
Mean R@1
2D-TAN [83]
8.9
3.2
0.9
5.5
10.0
3.8
1.2
8.0
10.3
4.2
1.5
8.6
VSLNet [81]
9.2
3.1
1.1
6.1
10.9
4.0
1.5
8.5
11.8
5.8
1.7
9.4
UMT [46]
9.7
3.5
1.2
6.5
11.2
5.4
1.7
8.7
12.0
6.1
1.6
9.4
STALE [57]
10.0
3.8
1.2
6.9
11.6
5.5
1.5
8.8
12.1
6.1
1.7
9.6
Ours
12.0
5.9
1.5
8.9
14.6
6.9
2.2
11.2
16.7 7.7 2.8 12.8
Table 2: Comparison of our method with prior methods at different input combinations for detour video retrieval (left) and detour window localization (right) on all metrics. Our method outperforms all the prior works for all input combinations.
4.2 Results at different input combinations
Tab. 2 contains an expanded version of Tab. 2 from the main paper for all metrics. We showed performance only on one metric for brevity. We see that for both the tasks, the previous source video context and the query context is useful for the model. It is also interesting to note that for state-of-the-art methods InternVideo [70] and CLIP [61], combining source video context directly with query features degrades the performance. This underscores the need for a smarter method to fuse the two contexts, as we show in our method.